M 361 54235 Fall ’98 T Th 12:30-2:00 RLM 5.124

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30pm or by appt. (Web page: http://www.ma.utexas.edu/˜ friedman)

Text and Syllabus: Basic Complex Analysis , 2nd ed. Marsden, Hoffman; Freeman Syllabus: Most of Chapters I-IV, more if time permits.

Prerequisites: M427K or M427L or consent of instructor.

Exams: We will have exams on 9/29 (class 10; Tues.), 11/3 (class 20; Tues.), 12/3 (Class 28; Thurs.) (No “Final exam”). All exam grades will be curved.

* Exam grades will be scaled (if the unscaled average is ≤ 75) to some value in the 75 to 80 range. MAKEUPS WILL BE GIVEN VERY RELUCTANTLY

Homework: Homework problems will be assigned each Tuesday in class and will be due the next Tuesday. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework together will be counted as a grade equivalent to one exam grade. (Homework assignments will be listed on my web page - see above.)

Final Grades: In computing a final grade, I will count the three tests and one total homework grade equally. (So there will be four grades to average.)